OpenStudy (anonymous):

If I have a fraction where the top and bottom have the same exponent, can I simplify? e.g. (n-1)^k / n^k

OpenStudy (anonymous):

yes, you can.

OpenStudy (anonymous):

So, \[((n-1)/n)^k\]?

OpenStudy (anonymous):

or further, \[(1 - (1/n))^k\] i guess?

OpenStudy (anonymous):

both OK.

OpenStudy (anonymous):

Cool, thanks. I guess this is because \[(2/3)^2 = 2/3 × 2/3 = 2^2/3^2\]

OpenStudy (anonymous):

it works even if k is a fraction or negative but you don't need to worry about it for now.